OpenStudy (anonymous):

How would I solve 1/x+1 = x/2

OpenStudy (anonymous):

I know to cross multiply and can get to x^2 + x = 2. stumped after that.

OpenStudy (anonymous):

\[\frac{1}{x+1}=\frac{x}{2}\iff x(x+1)=2\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

(x-1)(x+2) is a solution. so x = 1 and -2

OpenStudy (anonymous):

That can be factored. look into it

OpenStudy (anonymous):

yes you are right and i am wrong two solutions

OpenStudy (anonymous):

I guess I should ask do you understand the factorization @fiedster23 . I'm sure the quadratic would give you the same answer, but I haven't did it.

OpenStudy (anonymous):

yes and thank you. i got it now.